# Robot Navigation with Entity-Based Collision Avoidance using Deep Reinforcement Learning

### v1

@@ 6 Conclusion
In this study, we introduced a novel reinforcement learning approach, termed Entity-Based Collision Avoidance using Deep Reinforcement Learning (EB-CADRL), for the navigation of autonomous delivery robots in dynamic environments. Our primary focus was on enhancing safety and efficiency by leveraging entity-specific information to guide the robot’s interactions with different types of agents and obstacles. Our proposed reward function penalized the robot differently based on the type of entity involved in a collision and its proximity to entities, thereby ensuring more context-aware and safer navigation. Our experimental results demonstrate that EB-CADRL significantly outperforms traditional navigation methods and state-of-the-art techniques. Notably, our approach achieved a higher success rate in reaching the goal while maintaining a lower collision rate across various entity types. The model effectively balanced the trade-off between safety and efficiency, maintaining safe distances from entities without significantly increasing travel time. The ablation experiments further validated the importance of incorporating entity-specific information, as models with this feature consistently performed better in terms of reward, success rate, and overall safety metrics compared to those without it. This underscores the relevance and added value of our work in developing more sophisticated and context-aware navigation systems for autonomous robots. In conclusion, our proposed EB-CADRL approach represents a significant advancement in the field of autonomous robot navigation, offering a robust, safe, and efficient solution for navigating dynamic, crowded environments.
Appendix A Appendix
In this section we show figures of collision rate for each entity type on train and validation data.
Figure 5: Collision Rate by Entity Type on Train and Validation data.

### v2

@@ 6 Conclusion
In this study, we introduced Entity-Based Collision Avoidance using Deep Reinforcement Learning (EB-CADRL), a novel approach for autonomous robot navigation in environments with dynamic agents and static obstacles. To enhance safety and efficiency, EB-CADRL leverages an entity-specific reward function that ensures more context-aware navigation by penalizing the robot differently based on entity type. Our experimental results demonstrate that EB-CADRL significantly outperforms state-of-the-art techniques, achieving a high success rate with a lower collision rate across various entity types. The model also effectively balanced safety and efficiency, maintaining safe distances from entities without significantly increasing travel time. Furthermore, ablation experiments validated the critical role of incorporating entity-specific information, which consistently improved performance across reward, success, and safety metrics. This result underscores the value of our context-aware approach in developing more sophisticated autonomous navigation systems.
Future work could be devoted to enhancing the model’s robustness in environments with occlusions and complex agent behaviors, integrating additional entity types into the model, such as cars and motorcycles, exploring multi-robot coordination, and conducting real-world testing.